# LDP for invariant measures of the Allen-Cahn equation with inhomogeneous boundary conditions and multiplicative noise

### v1

@@ Abstract
We prove the validity of a small noise large deviation principle for the family of invariant measures $\{\mu_{\epsilon}\}_{\epsilon>0}$ associated to the one dimensional stochastic Allen-Cahn equation with inhomogeneous Dirichlet boundary conditions, perturbed by unbounded multiplicative noise. The main difficulty is that the system is not strongly dissipative. Using L. Simon’s convergence theorem, we show that the dynamics of the noiseless system converge in large time to the minimizer of the Ginzburg-Landau energy functional, which is unique due to the boundary condition. We obtain an estimate of the invariant measure on the bounded set in the Sobolev space $W^{k^{\star},p^{\star}}$, where $k^{\star}p^{\star}>1$, and $p^{\star}$ is large. As a corollary of the main result, we show that $\mu_{\epsilon}$ concentrates around the unique minimizer with such boundary conditions exponentially fast when $\epsilon$ is sufficiently small.
MSC2020 subject classifications: Primary 60H10, 60B10; secondary 37A50.
Keywords: large deviations principle; invariant measures; stochastic Allen-Cahn equation; quasi-potential
@@ 1 Introduction
In this paper, we study the following equation.
$\begin{cases}\partial_{t}u_{\epsilon}(t,\xi)=\Delta u_{\epsilon}(t,\xi)+u_{\epsilon}(t,\xi)-u_{\epsilon}(t,\xi)^{3}+\epsilon^{\frac{1}{2}}g(t,\xi,u_{\epsilon}(t,\xi))\partial_{t}W^{L}(t,\xi),\xi\in(-L,L),\\ u_{\epsilon}(0,\xi)=x(\xi),\\ u_{\epsilon}(t,\xi)=\mathrm{sgn}(\xi),\,|\xi|=L,\;t\geq 0,\end{cases}$ (1.1)
where $W^{L}(t,\xi):=\sum_{k\in\mathbb{N}}e_{K}^{L}(\xi)\beta_{k}(t)$, $\{e_{k}\}$ is a complete orthogonal basis in $H:=L^{2}(-L,L)$, and $\{\beta_{k}(t)\}$ is a sequence of independent Brownian motions. The mapping $g:[0,\infty)\times[-L,L]\times\mathbb{R}\rightarrow\mathbb{R}$ is measurable, and for almost all $t\geq 0$, the mapping $g(t,\cdot,\cdot):[-L,L]\times\mathbb{R}\rightarrow\mathcal{L}(\mathbb{R})$ is continuous. Moreover, for any $\theta,\vartheta\in\mathbb{R}$ and almost all $t\geq 0$,
@@
and denote by $S(t)$ the heat semigroup on (L,L)(-L,L) with zero boundary conditions at the endpoints. We denote by $\mathcal{E}$ the set of all $u\in C[-L,L]$ with u(L)=1,u(L)=1u(-L)=-1,\,u(L)=1 and by $E$ the space of all $u\in C[-L,L]$ with u(L)=u(L)=0u(-L)=u(L)=0. Note that $\mathcal{E}$ is not a linear space. Then, the mild solution of Eq.(1.1) is defined as $u_{\epsilon}(t):=\bar{u}_{\epsilon}(t)+\psi$. Here,
@@
$\begin{cases}\partial_{t}\bar{u}_{\epsilon}(t,\xi)=\Delta\bar{u}_{\epsilon}(t,\xi)+F(\bar{u}_{\epsilon}(t))(\xi)+\epsilon^{\frac{1}{2}}G(t,\bar{u}_{\epsilon}(t))\partial_{t}W^{L}(\xi),\xi\in(-L,L),\\ \bar{u}_{\epsilon}(0)=x-\psi,\\ \bar{u}_{\epsilon}(t,\xi)=0,\,|\xi|=L,\;t\geq 0,\end{cases}$ (1.3)
$G(t,z)\phi(\xi):=g(t,\xi,z(\xi)+\psi(\xi))\phi(\xi)$ and $F(z)(\xi):=-(z(\xi)+\psi(\xi))^{3}+(z(\xi)+\psi(\xi))$, $z\in E$. Note that for each $z$ satisfies the zero boundary condition, $F(z)$ also satisfies the zero boundary condition, so the mild solution of Eq.(1.3) is well defined.
According to [cerrai2003stochastic](see Theorem 5.3, Theorem 5.5, and Proposition 6.1), for any initial datum $x\in E$ and $T>0$ , there exists a unique mild solution $\bar{u}^{\epsilon}\in L^{p}(\Omega;C([0,T];E))$ with $p>1$. Moreover, there exists $c_{p}>0$ such that
@@
converges weakly to some measure $\bar{\mu}_{\epsilon}$, which is an invariant measure of Eq.(1.3) (see [cerrai2003stochastic], Theorem 6.2). We assume the following condition.
@@
The study of small noise LDP for invariant measures of SPDEs began in the work of Sowers [sowers1992large]. Cerrai and Röckner [cerrai2005large] considered the stochastic reaction-diffusion equation with a strongly dissipative nonlinear term on a multi-dimensional space. Brzeźniak and Cerrai [brzezniak2017large] considered the stochastic Navier-Stokes equation on $\mathbb{T}^{2}$ with additive noise. Wang [wang2024large] considered the stochastic reaction-diffusion equation on unbounded domains under strongly dissipative conditions. In all these cases, the systems have 0 as their global attractors in the underlying space, following a dissipative argument. In this paper, the dynamics are only weakly dissipative, so such arguments fail to work. The specific potential $V$ we consider in this paper has a double well with 2 critical points $\pm 1$, at which the potential attains its minimum. Due to the boundary condition of our problem, there exists a unique solution for the equation
@@
Moreover, the solution of the noiseless system $z^{x}_{0}(t)$ is the $L^{2}$ gradient flow of the energy functional
$\mathbf{E}_{L}(u):=\int_{-L}^{L}\left[\frac{|u^{\prime}(\xi)|^{2}}{2}+V(u(\xi))\right]\;d\xi,\;u(\pm L)=\pm 1.$
@@
By L. Simon’s convergence theorem, we show that for any $x\in E$, $z^{x}_{0}(t)$ converges to the minimizer of $\mathbf{E}_{L}$, which is the unique solution of (1.6), as time goes to infinity. The minimizer $m_{L}$ is the stationary profile with two pure phases of alloys, magnetization, or fluids coexisting at the two ends of a bounded interval. In these contexts, the boundary models the effects of two different materials, magnetization, or fluids applied to the endpoints.
@@
Note that in the case when the noise is space-time white noise, the invariant measure $\mu$ has an exponential density with respect to the Brownian bridge measure [reznikoff2005invariant]. By the exponential tilting argument, as an alternative version of the Varadhan Lemma, the LDP for the invariant measure will follow immediately (see Theorem III.17 of [hollander2000large]). In the case that $L(\epsilon)$ depends on $\epsilon$ and goes to infinity as $\epsilon\rightarrow 0$, the limiting behaviour of the invariant measure $\mu_{\epsilon}$ relies on how fast $L(\epsilon)$ diverges. Such problems are considered by Weber [weber2010sharp], Otto-Weber-Westdickenburg [otto2014invariant], Bertini-Stella-Buttà [bertini2008dobrushin], and Bertini-Buttà-Gesù [bertini2025asymptotics]. For all the works mentioned above, the noise must be space-time white noise. In the case of the multiplicative noise we consider in this paper, the density and tilted LDP argument break down. As far as we know, this is the first paper to consider the LDP for the invariant measure of semilinear parabolic equations with double well potential and multiplicative noise.
@@
The typical approach to proving the LDP for invariant measures of SPDEs with multiplicative noise involves a crucial step, which is to obtain an exponential estimate for the invariant measure $\bar{\mu}$ outside of the compact subset $B_{W^{k^{\star},p^{\star}}(-L,L)}(0,R)$ on the space of continuous functions $E$. The noise we consider is multiplicative space-time white noise. The intensity of the noise only needs to be nondegenerate and Lipschitz; thus, it is of linear growth. The exponential estimate for $\bar{\mu}$ is actually quite challenging in our case. In Cerrai-Paskal [cerrai2022large] and Wang [wang2024large], they choose the compact subset to be the bounded set of Sobolev space $H^{1}$, and the intersection between $H^{1}$ and the weighted $L^{2}$ space, respectively. Then, they obtained the desired exponential estimate by applying Ito’s formula on $H^{1}$. In our case, this approach does not work because of the non-dissipative nature of our system and the regularity of our noise. Following some ideas from our previous work [BAI2026111284], using the smoothing property of the heat semigroup, we transform the estimate for the $W^{k^{\star},p^{\star}}$ norm of the solution to SPDE (1.3) to that of the stochastic convolution:
@@
We choose to estimate the $W^{k^{\star},p^{\star}}$ norm of $\bar{u}^{0}_{\epsilon}$ for large $p^{\star}$ such that $p^{\star}k^{\star}>1$ to ensure that the Sobolev space $W^{k^{\star},p^{\star}}$ is compactly embedded into $E$. The advantage is to allow relatively small $k^{\star}<\frac{1}{2}$. Then, the $W^{k^{\star},p^{\star}}$ estimate of $\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})$ can be obtained by the factorization method. This estimate is divided into two cases: probability conditioned on the $\bar{u}^{0}_{\epsilon}(t)$ being bounded and unbounded under the $E$ norm, respectively. When $\bar{u}^{0}_{\epsilon}$(t) is bounded; the estimate is pinned to the case when $g$ is bounded above; then we use the LDP technique to obtain the exponential estimate for the conditional probability. The probability that $\bar{u}^{0}_{\epsilon}(t)$ is outside a bounded subset of $E$ is obtained from its exponential expectation through Taylor expansion and moment estimates. The moment estimation method was suggested by Cerrai [cerrai2003stochastic]. With the exponential tightness result of $\mu_{\epsilon}$ and the fact that the noiseless dynamics have a unique global attractor, we are able to prove the LDP for $\mu_{\epsilon}$. Moreover, as a corollary, it is shown that $\mu_{\epsilon}$ concentrates around $m_{L}$, which is the non-trivial minimizer of $\mathbf{E}_{L}$, exponentially fast.
@@ 2 LDP for dynamics
The large deviations result for the dynamics of Eq.(1.3) has been well studied by Cerrai and Röckner in [cerrai2004large]. We formulate their results (see [cerrai2004large], Theorems 6.2 and 6.3) to obtain the Freidlin-Wentzell uniform large deviations principle (FWULDP) for the dynamics of Eq.(1.3):
@@
Let $\mathcal{M}$ be a Banach space and $D$ be some non-empty set. Suppose that for each $x\in D$, $\{\mu^{x}_{\epsilon}\}_{\epsilon>0}$ is a family of probability measures on $\mathcal{M}$ and an action function $I^{x}:\mathcal{M}\rightarrow[0,+\infty]$ is a good rate function. The family $\{\mu^{x}_{\epsilon}\}_{\epsilon>0}$ is said to satisfy a Dembo-Zeitouni uniform large deviation principle (DZULDP) in $\mathcal{M}$, with rate function $I^{x}$, uniformly with respect to $x\in D$, if the following hold.
@@
Since we have the FWULDP, according to Theorem 2.7 of [salins2019equivalences], it suffices the show that for any $x_{n}\rightarrow x\in E$,
@@
This is implied by the continuity of the skeleton function with respect to initial conditions. To see this, first since $F$ is locally Lipschitz continuous on $E$, there exists $C_{R}>0$ such that
@@
for any $x,y\in B_{R}(E)$. Moreover, due to Theorem 4.1 of [cerrai2004large], for $x,y\in E$ and $\phi,\varphi\in L^{2}(0,T;H)$, there exists $R$ such that
@@
By Grönwall inequality, this implies
@@
Therefore, by taking $\phi=\varphi$
$\sup_{z\in K_{T}^{x}}d_{C([0,T];E)}(z,K_{T}^{x_{n}}(r))\leq c(T,\|x\|_{E},r)\|x-x_{n}\|_{E},$

### v2

# Large deviations for invariant measures of the stochastic Allen-Cahn equation with inhomogeneous boundary conditions and multiplicative noise

@@ Abstract
We prove the validity of a small noise large deviation principle for the family of invariant measures $\{\mu_{\epsilon}\}_{\epsilon>0}$ associated to the one dimensional stochastic Allen-Cahn equation with inhomogeneous Dirichlet boundary conditions, perturbed by unbounded multiplicative noise. The novelty is that the system is not strongly dissipative, and the noise intensity can be of linear growth while the drift term is of polynomial growth. Using L. Simon’s convergence theorem, we show that the dynamics of the noiseless system converge in large time to a minimiser of the Ginzburg-Landau energy functional, which is unique due to the boundary condition. We obtain an estimate of the invariant measure on a bounded set in the Sobolev space $W^{k^{\star},p^{\star}}$, where $k^{\star}p^{\star}>1$, and $p^{\star}$ is large, which is compactly embedded in the continuous function space. As a consequence of the LDP result, we show that $\mu_{\epsilon}$ concentrates around the unique minimiser with such boundary conditions exponentially fast when $\epsilon$ is sufficiently small.
MSC2020 subject classifications: Primary 60H15, 60F10, 35R60; secondary 37A50, 37L55.
Keywords: large deviation principle; invariant measures; stochastic Allen-Cahn equation; unbounded multiplicative noise; inhomogeneous boundary conditions.
@@ 1 Introduction
In this paper, we study the following equation
$\begin{cases}\partial_{t}u_{\epsilon}(t,\xi)=\Delta u_{\epsilon}(t,\xi)+u_{\epsilon}(t,\xi)-u_{\epsilon}(t,\xi)^{3}+\epsilon^{\frac{1}{2}}g(t,\xi,u_{\epsilon}(t,\xi))\partial_{t}W(t,\xi),\xi\in(-L,L),\\ u_{\epsilon}(0,\xi)=x(\xi),\\ u_{\epsilon}(t,\xi)=\mathrm{sgn}(\xi),\,|\xi|=L,\;t\geq 0,\end{cases}$ (1.1)
where $W(t,\xi):=\sum_{k\in\mathbb{N}}e_{k}(\xi)\beta_{k}(t)$, $\{e_{k}\}$ is a complete orthogonal basis in $H:=L^{2}(-L,L)$, and $\{\beta_{k}(t)\}$ is a sequence of independent real valued Brownian motions on the probability space $(\Omega,\mathcal{F},\mathcal{F}_{t},\mathbb{P})$. Assume that the mapping $g:[0,\infty)\times[-L,L]\times\mathbb{R}\rightarrow\mathbb{R}$, is measurable, and for almost all $t\geq 0$, $g(t,\cdot,\cdot):[-L,L]\times\mathbb{R}\rightarrow\mathcal{L}(\mathbb{R})$, is continuous. Moreover, for any $\theta,\vartheta\in\mathbb{R}$ and almost all $t\geq 0$,
@@
and denote by $S(t)$ the heat semigroup on (L,L)(-L,L) with zero boundary conditions at the endpoints. We denote by $\mathcal{E}$ the set of all $u\in C[-L,L]$ with u(L)=1,u(L)=1u(-L)=-1,\,u(L)=1 and by $E$ the space of all $u\in C[-L,L]$ with u(L)=u(L)=0u(-L)=u(L)=0. Note that $\mathcal{E}$ is not a linear space. Then, the mild solution of Eq.(1.1) is defined as $u_{\epsilon}(t):=\bar{u}_{\epsilon}(t)+\psi$, and $\bar{u}_{\epsilon}$ satisfies
@@
$\begin{cases}\partial_{t}\bar{u}_{\epsilon}(t,\xi)=\Delta\bar{u}_{\epsilon}(t,\xi)+F(\bar{u}_{\epsilon}(t))(\xi)+\epsilon^{\frac{1}{2}}G(t,\bar{u}_{\epsilon}(t))\partial_{t}W(\xi),\xi\in(-L,L),\\ \bar{u}_{\epsilon}(0)=x-\psi,\\ \bar{u}_{\epsilon}(t,\xi)=0,\,|\xi|=L,\;t\geq 0.\end{cases}$ (1.3)
Here, $G(t,z)\phi(\xi):=g(t,\xi,z(\xi)+\psi(\xi))\phi(\xi)$ and $F(z)(\xi):=-(z(\xi)+\psi(\xi))^{3}+(z(\xi)+\psi(\xi))$, $z\in E$. Note that for each $z$ that satisfies the zero boundary condition, $F(z)$ also satisfies the zero boundary condition, so the mild solution of Eq.(1.3) is well defined.
According to [cerrai2003stochastic] (see Theorem 5.3, Theorem 5.5, and Proposition 6.1), for any initial datum $x\in E$ and $T>0$ , there exists a unique mild solution $\bar{u}_{\epsilon}^{x}\in L^{p}(\Omega;C([0,T];E))$ with $p>1$. Moreover, there exists $c_{p}>0$ such that
@@
converges weakly to some measure $\bar{\mu}_{\epsilon}$, which is an invariant measure of Eq.(1.3) (see [cerrai2003stochastic], Theorem 6.2). We assume the following condition,
@@
The study of small noise LDP for invariant measures of SPDEs began in the work of Sowers [sowers1992large]. Cerrai and Röckner considered the stochastic reaction-diffusion equation with a strongly dissipative nonlinear term in a multi-dimensional space ([cerrai2005large]). Brzeźniak and Cerrai studied the stochastic Navier-Stokes equation on $\mathbb{T}^{2}$ with additive noise ([brzezniak2017large]). Wang investgated the stochastic reaction-diffusion equation on an unbounded domain under strongly dissipative conditions ([wang2024large]). Bai, Feng and Zhao studied the stochastic Burgers equation on the $(0,1)$ domain ([BAI2026111284]). In all these cases, the systems have 0 as their global attractor in the underlying space, following a dissipative argument. In this paper, the dynamics are only weakly dissipative and the existing arguments fail to work for our situation. The specific potential $V(u)=\frac{1}{4}(u^{2}-1)^{2}$ we consider in this paper has double wells with 2 critical points $\pm 1$ where the potential attains its minimum. Due to the boundary condition of our problem, there exists a unique solution for the following equation
@@
Moreover, the solution of the noiseless system $u_{0}(t)$ (i.e. $\epsilon=0$ in equation (1.1)) is the $L^{2}$ gradient flow of the energy functional
$\mathbf{E}_{L}(u):=\int_{-L}^{L}\left[\frac{|u^{\prime}(\xi)|^{2}}{2}+V(u(\xi))\right]\;d\xi,\;u(\pm L)=\pm 1.$ (1.7)
@@
By using L. Simon’s convergence theorem, we show that for any $x\in E$, $u_{0}(t)$ converges to the minimizer of $\mathbf{E}_{L}$, which is the unique solution of (1.6), as time goes to infinity. The minimizer $m_{L}$ is the stationary profile with two pure phases of alloys, magnetization, or fluids coexisting at the two ends of a bounded interval. In these contexts, the boundary models the effects of two different materials, magnetization, or fluids applied to the endpoints.
@@
Note that in the case where the noise is space-time white noise, the invariant measure $\mu$ has an exponential density with respect to the Brownian bridge measure [reznikoff2005invariant]. By the exponential tilting argument, as an alternative version of the Varadhan Lemma, the LDP for the invariant measure will follow immediately (see Theorem III.17 of [hollander2000large]). In the case where $L(\epsilon)$ depends on $\epsilon$ and goes to infinity as $\epsilon\rightarrow 0$, the limiting behaviour of the invariant measure $\mu_{\epsilon}$ depends on the rate at which $L(\epsilon)$ diverges. Such problems are considered by Weber [weber2010sharp], Otto-Weber-Westdickenburg [otto2014invariant], Bertini-Stella-Buttà [bertini2008dobrushin], and Bertini-Buttà-Gesù [bertini2025asymptotics]. For all the works mentioned above, the noise must be additive space-time white noise. In the case of the multiplicative noise we consider here, the density and tilted LDP arguments break down. As far as we know, this is the first paper to consider the LDP for the invariant measure of semilinear parabolic equations with a double-well potential and multiplicative noise.
@@
The typical approach to proving the LDP for invariant measures of SPDEs with multiplicative noise involves a crucial step to obtain an exponential estimate for the invariant measure $\bar{\mu}$ outside of the compact subset $B_{W^{k^{\star},p^{\star}}(-L,L)}(0,R)$ of the space of continuous functions $E$. The noise we consider is multiplicative space-time white noise. The intensity of the noise only needs to be Lipschitz, so it is of linear growth. Note that we impose a nondegeneracy condition on the noise in order to analyze the relationship between the energy functional and the rate functional. It is worth mentioning that in Cerrai-Röckner [cerrai2005large], when the nonlinear term is cubic but still dissipative, the noise intensity cannot have linear growth. The exponential estimate for $\bar{\mu}$ is actually quite challenging in our case. In [cerrai2022large], the authors chose the compact subset to be a bounded set of the Sobolev space $H^{1}$, while Wang [wang2024large] dealt with trace-class additive noise in $H^{1}(\mathbb{R})$, and chose the compact subset as the intersection of bounded sets in $H^{1}$ and a weighted $L^{2}$ space. This was followed by the desired exponential estimate when applying Ito’s formula on $H^{1}$. However, this approach does not work in our case because of the non-dissipative nature of our system and the lower regularity and growth rate of the intensity of the noise.
Motivated by some ideas from the previous work [BAI2026111284], and using the smoothing property of the heat semigroup, we transform the estimate for the $W^{k^{\star},p^{\star}}$ norm of the solution of the SPDE (1.3) to that of the following stochastic convolution:
@@
We choose to estimate the $W^{k^{\star},p^{\star}}$ norm of $\bar{u}^{0}_{\epsilon}$ for large $p^{\star}$ such that $p^{\star}k^{\star}>1$ to ensure that the Sobolev space $W^{k^{\star},p^{\star}}$ is compactly embedded in the space $E$. Then, the $W^{k^{\star},p^{\star}}$ estimate of $\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})$ can be obtained by the factorisation method under the condition that the noise intensity is bounded. When $g$ is unbounded, the estimate is split into two terms: probability conditioned on the set $\bar{u}^{0}_{\epsilon}(t)$ being bounded and unbounded under the $E$ norm, respectively. Conditioned on $\bar{u}^{0}_{\epsilon}(t)$ being bounded, the estimate is pinned to the case when $g$ is bounded from above. Then we use the LDP technique to obtain the exponential estimate for the conditional probability. The probability of $\bar{u}^{0}_{\epsilon}(t)$ being outside a bounded subset of $E$ is obtained from the variational principle, which was also used in Budhiraja-Dupuis-Maroulas [budhiraja2008large]. The advantage of this approach is that we avoid estimating the exponential expectation of $\bar{u}_{\epsilon}^{0}$ directly, which allows the noise intensity to grow linearly. We believe that this approach can be used to study large deviations for invariant measures of other SPDEs. With the exponential tightness result of $\mu_{\epsilon}$ and the fact that the corresponding noiseless dynamics has a unique global attractor, we are able to prove the LDP for $\mu_{\epsilon}$. Moreover, as a corollary, it is shown that $\mu_{\epsilon}$ concentrates around $m_{L}$, which is the non-trivial minimiser of $\mathbf{E}_{L}$, exponentially fast in $\epsilon$.
@@ 2 LDP for dynamics
The large deviation result for the dynamics of Eq.(1.3) has been well studied by Cerrai and Röckner in [cerrai2004large]. We formulate their results (see [cerrai2004large], Theorems 6.2 and 6.3) to obtain the Freidlin-Wentzell uniform large deviations principle (FWULDP) for the dynamics of Eq.(1.3):
@@
Let $\mathcal{M}$ be a Banach space, and $D\subset\mathcal{M}$ be some non-empty set. Suppose that for each $x\in D$, $\{\mu^{x}_{\epsilon}\}_{\epsilon>0}$ is a family of probability measures on $\mathcal{M}$ and the action function $I^{x}:\mathcal{M}\rightarrow[0,+\infty]$ is a good rate function. The family $\{\mu^{x}_{\epsilon}\}_{\epsilon>0}$ is said to satisfy a Dembo-Zeitouni uniform large deviation principle (DZULDP) in $\mathcal{M}$, with rate function $I^{x}$, uniformly with respect to $x\in D$, if the following estimates hold:
@@
Since we have FWULDP, according to Theorem 2.7 of [salins2019equivalences], it suffices to show that for any $x_{n}\rightarrow x\in E$,
@@
This is implied by the continuity of the skeleton function with respect to the initial conditions. To see this, first since $F$ is locally Lipschitz continuous on $E$, there exists $C_{R}>0$ such that
@@
for any $x,y\in B_{R}(E)$. Moreover, according to Theorem 4.1 of [cerrai2004large], for $x,y\in E$ and $\phi,\varphi\in L^{2}(0,T;H)$, there exists $R$ such that
@@
By the Grönwall inequality, this implies
@@
Therefore, for any $z\in K_{T}^{x}$, there exists $\phi$ such that $z=z_{\phi}^{x}$ and $\frac{1}{2}\|\phi\|_{L^{2}(0,T;H)}\leq r$. Then, we have
$d_{C([0,T];E)}(z,K_{T}^{x_{n}}(r))\leq d_{C([0,T];E)}(z_{\phi}^{x},z_{\phi}^{x_{n}})\leq c(T,\|x\|_{E},r)\|x-x_{n}\|_{E},$